Fermi National Laboratory
 

Operation Summary for May 10, 2010 - May 12, 2010

 

Monday May 10

The day shift began with Operations monitoring the Tevatron (TeV) store 7804 (100.88E30) and the Recycler stash (215.6E10), with the Antiproton Source (Pbar) stacking (17.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations reported that the Booster Notcherr and NuMI kicker power supplies caused some short downtimes.

During the evening shift, Operations terminated store 7804 at 4:15 PM, disabled the MTest beam at 4:24 PM, put the TeV into HEP shot setup, and established store 7805 at 5:45 PM with an average initial luminosity (AIL) of 377.8E30.

Tuesday May 11

The midnight shift began with Operations monitoring store 7805 (138.94E30) and the stash (188.4E10), with Pbar stacking (19.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 2:15 AM, the NuMI Lambertson power supply tripped off on a ground fault. Operations called in power supply personnel to investigate, which lasted throughout the shift. Operations established beam to MTest at 4:30 AM. Operators had the NuMI Target Hall and Absorber enclosures prepared for access by 7 AM.

During the day shift, Operations terminated store 7805 at 8:34 AM, put the TeV shot setup, and established store 7806 at 10:33 AM with an AIL of 361.8E30. The NuMI access continued. Power supply personnel completed the NuMI Lambertson repairs. A MI RF clock glitch caused an hour of downtime.

During the evening shift, NuMI personnel completed their access at 5:15 PM. Operations disabled the MTest beam at 6:05 PM. Operators established beam to NuMI at 6:30 PM.

Wednesday May 12

The midnight shift began with Operations monitoring store 7806 (78.44E30) and the stash (317.9E10), with Pbar stacking (15.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7806 at 2:58 AM, put the TeV into shot setup, and established store 7807 at 4:10 AM with an AIL of 335.1E30. Operations disabled the NuMI beam at 5 AM for a NuMI access that started at 7:39 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/12/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory